OpenStudy (anonymous):

Write the ratio: 2.2 to 10 as a fraction in lowest terms

OpenStudy (anonymous):

11/50

OpenStudy (anonymous):

how do you do it?

OpenStudy (anonymous):

\[\frac{2.2*10}{10*10}=\frac{22}{100}=\frac{2*11}{2*50}=\frac{11}{50}\]

OpenStudy (anonymous):

thank you